OpenStudy (anonymous):

convert directly to base 1O: 2ABBC base 16

ganeshie8 (ganeshie8):

2ABBC \(2 \times 16^4 + A \times 16^3 + B \times 16^2 + B \times 16^1 + C \times 16^0\)